solve using the Multiplication principle first then add
X=3Y, 5X+14=Y

Answer by checkley75(3666)

MULTIPLY THE FIRST EQUATION BY -5 THUS -5X=-15Y NOW ADD:
-5X=-15Y
5X+14=Y
-----------------------
14=-14Y
Y=14/-14
Y=-1 ANSWER.
X=3*-1
X=-3 ANSWER.
PROOF
5*-3+14=-1
-15+14=-1
-1=-1